OpenStudy (anonymous):

how would i use the x,y, and r definitions of sinx and cosx to prove the identity (cosθ)/1-sinθ=(1+sinθ)/cosθ

OpenStudy (anonymous):

sinx=y/r cosx=x/r

OpenStudy (anonymous):

this is what i did \[(x/r)/(1-y/r)= (1+y/r)/(x/r)\]\[(x/r)-(y/r)=(x/r)+(y+r)\]\[(x-y)/r=(x+y)/r\] ... :S

OpenStudy (anonymous):

cross multiply 1º line: \[(x/r)^{2}=1-(y/r)^{2}----->\cos ^{2}x=1-\sin ^{2}x\]

OpenStudy (anonymous):

so i have to do (x/r)^2/(1-y/r)^2 = (1+y/r)^2/(x/r)^2 and then work from there? :S ....

OpenStudy (anonymous):

no

OpenStudy (anonymous):

from what you did, take first line and cross multiply it

OpenStudy (anonymous):

(x/r)(x/r)=(1+y/r)(1-y/r)

OpenStudy (anonymous):

so i move the x/r to the left side and the 1-y/r to the right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yep. since (x/r)^2 = 1-(y/r)^2 or cos^2x=1-sin^2x

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

yw